state college has 620 stundent. there are 20 more women than men . how many women were there"
Answer by drk(1908)

Let M = men and W = women.
Now
M + W = 620
but we have
W = 20 + M.
So, we get
M + (20+M) = 620
simplifying, we get
2M + 20 = 620
2M = 600
M = 300
W = 320